OpenStudy (anonymous):

integral of cos (x/2)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Hello, DS! What is the integral of plain cos x? Bet you know that.

OpenStudy (anonymous):

sin x

OpenStudy (anonymous):

my mind's blank right now...

OpenStudy (mathmale):

I know the feeling! :) The integral of cos x is indeed sin x (plus C). Now, back to your post. What is the integral of cos (x/2)?

OpenStudy (mathmale):

You might want to try using the substitution u=x/2.

OpenStudy (anonymous):

no clue... is it sin x/2 ?

OpenStudy (mathmale):

Almost. Let u=x/2. Then (du/dx)=dx/2. Can you agree with that?

OpenStudy (anonymous):

sin x^2 /4

OpenStudy (mathmale):

In the long run it'd be more productive to show all your work. That way I can give you more meaningful feedback than "right" or "wrong." If we let u=x/2, then du/dx = 1/2, so that du = dx/2, or 2du = dx. Then your integral of cos x/2 with respect to x becomes \[\int\limits_{}^{}\cos u (2du)\rightarrow 2\int\limits_{}^{}\cos u du=??\]

OpenStudy (anonymous):

cos x/4

OpenStudy (mathmale):

\[\int\limits\limits_{}^{}\cos u (2du)\rightarrow 2\int\limits\limits_{}^{}\cos u du=2 \sin u + C\] Please make sure you're comfortable with this before we proceed further.

OpenStudy (anonymous):

i meant

OpenStudy (anonymous):